OpenStudy (anonymous):

This is the only question on the entire quiz I don't get. Will fan + medal, please help!!! A student is trying to solve the system of two equations given below: Equation P: x + y = 8 Equation Q: 2x + 5y = 24 Which of the following steps can be used to eliminate the x term? 2(x + y = 8) −2(x + y = 8) −1(2x + 5y = 24) −2(2x + 5y = 24)

OpenStudy (amistre64):

what does it mean to eliminate something?

OpenStudy (anonymous):

@amistre64 To get rid of?

OpenStudy (amistre64):

yeah, or in math terms .. make it go to zero

OpenStudy (amistre64):

so, what are our x terms in this case?

OpenStudy (anonymous):

@amistre64 2x and x?

OpenStudy (amistre64):

youre doing fine, yes now, what would you propose we do to eliminate them? what would we multiply one by such that when we add them together we get zero?

OpenStudy (anonymous):

I wanna say 8 because if we're multiplying 1, 1 x 8 is 8. But at the same time I feel like that's wrong?

OpenStudy (amistre64):

hmm, 8 isnt in our options so your feeling is spot on ... let "n" be some number such that: 2x + nx = 0, fpr all values of x (2+n)x = 0 any more ideas?

OpenStudy (anonymous):

Oh! I was looking at the first equation. Would it be 5 then?

OpenStudy (anonymous):

Or 2? Since we're trying to get rid of the x...(I'm sorry I know I'm really slow in math)

OpenStudy (amistre64):

2 is closer :) lets try it out (2+2)x = 0 4x = 0, this is not true for all x when is 2+n = 0 ?

OpenStudy (anonymous):

2+n = 0 when n = -2, right?

OpenStudy (amistre64):

that better, yes so what this tells us is that we want to multiply by a -2 the last thing to determine is which equation should we multiply by -2?

OpenStudy (anonymous):

Equation Q because if we solve that one first we can just replace the x and y variables in question P?

OpenStudy (amistre64):

well, equation Q already has 2x in it ... so we really want to leave it as is. this leaves us with equation P. -2(x+...) = -2x + ... now when we combine (add) -2P and Q we get -2x + ... 2x + ... --------- 0x + ... and our x term has been eliminated.

OpenStudy (amistre64):

so, which option should we pick now?

OpenStudy (anonymous):

B?

OpenStudy (amistre64):

hmm, your post does not letter them, but i would agree with the second option yes

OpenStudy (anonymous):

Oh, sorry! Also do you mind if I save this? Like your responses? You helped me a lot

OpenStudy (amistre64):

save what you want :)

OpenStudy (anonymous):

Thank you for helping me!!!

OpenStudy (amistre64):

youre welcome, and good luck :)